OpenStudy (anonymous):

What is the difference? nine fourths – one seventh (1 point) negative three eighths fifty nine twenty eighths negative eight thirds sixty seven twenty eighths

OpenStudy (ghazi):

\[\frac{ 9 }{ 4 }-\frac{ 1 }{ 7 }=\frac{ 9*7-4*1 }{ 28 }=\frac{ 63-4 }{ 28 }\] just take the LCM of denominators